OpenStudy (eric_d):

Chap 5 (EXE 5.4) 5)a)

OpenStudy (eric_d):

@rational

OpenStudy (eric_d):

http://prntscr.com/4l7kql

OpenStudy (rational):

You're given : gradient of " `xy = c^2`" at a point `(ct, c/t)` is `-1/t^2`

OpenStudy (rational):

5a : xy = 9

OpenStudy (rational):

compare this with the given form `c = ?`

OpenStudy (eric_d):

3 ?

OpenStudy (rational):

yes what about t ?

OpenStudy (eric_d):

I'm nt sure Actually, t refers to parameter ?

OpenStudy (rational):

t is t itself, so the gradient would be simply `-1/t^2`

OpenStudy (rational):

lets do 5b

OpenStudy (eric_d):

ok

OpenStudy (rational):

5b : xy = 4

OpenStudy (rational):

`c = ?`

OpenStudy (eric_d):

3 and 4/3 ?

OpenStudy (rational):

c = ?

OpenStudy (eric_d):

4^2 ?

OpenStudy (rational):

xy = 4 xy = 2^2 c = 2

OpenStudy (rational):

next look at given point : (3, 4/3) compare this with (ct, c/t)

OpenStudy (rational):

ct = 3 t = 3/c = 3/2

OpenStudy (rational):

yes ?

OpenStudy (eric_d):

ok

OpenStudy (rational):

so the gradient would be ?

OpenStudy (rational):

gradient = -1/t^2 = ?

OpenStudy (eric_d):

I really don't know..

OpenStudy (rational):

plugin the value of t here

OpenStudy (rational):

t = 3/2

OpenStudy (eric_d):

gradient=